### v2

Abstract
For a chain of cycles $\Gamma$ we prove that $\operatorname{Cliff}(\Gamma)=\operatorname{gon}(\Gamma)-2$.
2000 Mathematics Subject Classification.
05C25, 14T15
Keywords.
Clifford index, gonality, tableaux, metric graphs
1 Introduction
During the recent decades, a theory of divisors on metric graphs is developed having lots of properties similar to those on smooth projective curves. For the definitions of this theory and the motivations to develop such a theory we refer to [ref1].
Using this theory one can define, on a metric graph $\Gamma$, the gonality $\operatorname{gon}(\Gamma)$ (see Definition2) and the Clifford Index $\operatorname{Cliff}(\Gamma)$ (see Definition 4) in the same way as in the case of a smooth curve $C$. As in the case of smooth curves the inequality $\operatorname{Cliff}(\Gamma)\leq\operatorname{gon}(\Gamma)-2$ holds by definition. In the case of a smooth curve $C$ the inequality $\operatorname{gon}(C)-3\leq\operatorname{Cliff}(C)$ is proved in [ref2]. Whether or not this inequality also holds for metric graphs is mentioned as an unknown fact in [ref3]. The importance of this inequality comes from the fact that $\operatorname{gon}(\Gamma)$ is defined using divisors of rank 1 on $\Gamma$, while $\operatorname{Cliff}(\Gamma)$ gives information on divisors on $\Gamma$ of any prescribed rank.
In this paper we consider this question for chains of cycles and we prove the following result.
Theorem 1.
If $\Gamma$ is a chain of cycles of genus $g\geq 3$ then $\operatorname{Cliff}(\Gamma)=\operatorname{gon}(\Gamma)-2$
Chains of cycles were considered for the first time in [ref4] where the authors used those metric graphs to obtain a tropical proof of the Brill-Noether Theorem. Later on those metric graphs are used to prove many results in Brill-Noether Theory (see some applications in [ref1], see e.g. [ref5]).
Although the statement in the theorem is stronger than the result for curves in [ref2], this stronger statement does not hold for arbitrary metric graphs. As an example, if $\Gamma$ is a metric graph having the complete graph $K_{n}$ as an underlying finite graph, then $\operatorname{gon}(\Gamma)=n-1$ while $\operatorname{Cliff}(\Gamma)=n-4$ (see [ref6], including more results concerning divisors on such graphs).
In the proof of the theorem we rely on the results from [ref7] concerning divisor theory on chains of cycles. In particular from this theory the theorem is translated into a statement on so-called $\underline{m}$-displacement tableaux (see Theorem 2).
In the second section we give a short overview on the results of [ref7] and we make the statement on $\underline{m}$-displacement tableaux. In the third section we give a proof of that statement.
2 Some results on divisor theory on chains of cycles
Let $\Gamma$ be a metric graph of genus $g\geq 2$.
Definition 1.
The rank $r=\operatorname{rk}(D)$ of a divisor $D$ on $\Gamma$ is the largest integer $n$ such that for each effective divisor $E$ of degree $n$ on $\Gamma$ there exists an effective divisor $D^{\prime}$ on $\Gamma$ equivalent to $D$ containing $E$. (In case $D$ is not equivalent to an effective divisor then we put $\operatorname{rk}(D)=-1$.)
Definition 2.
The gonality $n=\operatorname{gon}(\Gamma)$ of $\Gamma$ is the smallest integer $m$ such that there exists a divisor $D$ on $\Gamma$ of degree $m$ such that $\operatorname{rk}(D)\geq 1$.
On a metric graph $\Gamma$ of genus $g$ we have a canonical divisor $K_{\Gamma}$ and there is a Riemann-Roch Theorem as in the case of curves. Divisors having rank at most 0 are not interesting, so we are only interested in divisors $D$ satisfying $\operatorname{rk}(D)\geq 1$ and $\operatorname{rk}(K_{\Gamma}-D)\geq 1$. Because of the Riemann-Roch Theorem this second condition is equivalent to $\operatorname{rk}(D)-d+g\geq 2$.
Definition 3.
The Clifford index of a divisor $D$ on $\Gamma$ is $\operatorname{Cliff}(D)=\deg(D)-2\operatorname{rk}(D)$.
The Clifford inequality states $\operatorname{Cliff}(D)\geq 0$ if $\operatorname{rk}(D)\geq 1$ and $\operatorname{rk}(D)-d+g\geq 2$.
Definition 4.
The Clifford index $c=\operatorname{Cliff}(\Gamma)$ is the smallest integer $m$ such that there exists a divisor $D$ with $\operatorname{rk}(D)\geq 1$ and $\operatorname{rk}(D)-d+g\geq 2$ with $\operatorname{Cliff}(D)=m$.
From now on $\Gamma$ is a chain of cycles of genus $g$. Such a chain is constructed as follows. Take $g$ graphs $C_{1},\cdots,C_{g}$ each of them isometric to a circle (those are the cycles). On each cycle $C_{i}$ we choose two different points $v_{i}$ and $w_{i}$. For $11\leq i\leq g-$ 1 the points $w_{i}$ and $v_{i+1}$ are connected by a straight line segment. We use the notation $(C_{i},v_{i},w_{i})$ throughout the paper.
Figure 1: a chain of cycles
The length of those segments does not influence the divisor theory on $\Gamma$. In particular it can be taken equal to zero. Taking it nonzero is handsome to distinguish e.g. the points $w_{1}$ on $C_{1}$ and $v_{2}$ on $C_{2}$. In order to prove Theorem 1 it is enough to prove the following statement on such a graph $\Gamma$. If $D$ is a divisor of degree $d$ and rank at least $r$ on $\Gamma$ with $r\geq 1$ and $r-d+g\geq 2$ then there exists a divisor $D^{\prime}$ of degree $d-2r+2$ and rank at least 1 on $\Gamma$.
We now state some definitions and results from [ref7] on chains of cycles $\Gamma$. Such a chain of cycles $\Gamma$ has torsion profile $\underline{m}=(m_{2},\cdots,m_{g})$ consisting of $g-1$ non-negative integers defined as follows.
Definition 5.
Let $l_{i}$ be the length of the cycle $C_{i}$ and let $l(v_{i},w_{i})$ be the length of the clockwise arc on $C_{i}$ from $v_{i}$ to $w_{i}$. We put $m_{i}=0$ if $l_{i}$ is an irrational multiple of $l(v_{i},w_{i})$. Otherwise $m_{i}$ is the minimal positive integer such that $m_{i}l(v_{i},w_{i})$ is an integer multiple of $l_{i}$.
Definition 6.
For positive integers $m$ and $n$ we write $[m\times n]$ to denote the set $\{1,\cdots,m\}\times\{1,\cdots,n\}$. An $\underline{m}$-displacement tableau on $[m\times n]$ of genus $g$ is a function $t:[m\times n]\rightarrow\{1,\cdots,g\}$ such that
  • $t$ is strictly increasing if one of the two coordinates is fixed.
  • if $t(x,y)=t(x^{\prime},y^{\prime})$ then $x-y\equiv x^{\prime}-y^{\prime}\mod m_{t(x,y)}$.
From the results in [ref7] we obtain
Proposition 1.
Given $r\in\mathbb{Z}_{\geq 1}$ and $d\in\mathbb{Z}$ with $r-d+g\geq 2$. There exists a divisor $D$ on $\Gamma$ of degree $d$ and rank at least $r$ if and only if there exists an $\underline{m}$-displacement tableau $t:[(g-d+r)\times(r+1)]\rightarrow\{1,\cdots,g\}$.
So in order to prove Theorem 1 we need to prove
Theorem 2.
Let $r\in\mathbb{Z}_{\geq 1}$ and $d\in\mathbb{Z}$ with $r-d+g\geq 2$. If there exists an $\underline{m}$-displacement tableau $t:[(g-d+r)\times(r+1)]\rightarrow\{1,\cdots,g\}$ then there exists an $\underline{m}$-displacement tableau $t^{\prime}:[(g-d+2r-1)\times 2]\rightarrow\{1,\cdots,g\}$.
3 Proof of the theorem
In case $r=1$ there is nothing to prove so we are going to assume $r\geq 2$ and we assume the statement of Theorem 2 holds in case $r$ is replaced by any smaller integer.
Lemma 1.
The statement of Theorem 2 holds for $d\geq g$.
Proof.
Assume $d\geq g$. Then $K_{\Gamma}-D$ is a divisor of degree $d^{\prime}=2g-2-d$ and rank at least equal to $r^{\prime}=r+g-d-1$. Since $g-d+r\geq 2$ we have $r^{\prime}\geq 1$ while $g-d^{\prime}+r^{\prime}=r+1\geq 2$. Moreover $r^{\prime}\leq r-1$, so we can apply the induction hypothesis and conclude that there is an $\underline{m}$-displacement tableau on $[(g-d^{\prime}+2r^{\prime}-1)\times 2]=[(g-d+2r-1)\times 2]$. ∎
So from now on in the proof we assume $d\leq g-1$.
In the proof we are going to use tableaux associated to ”smaller” chains of cycles. We use the following terminology. If $11\leq k\leq g-$ 1 then by $C_{1}\cup\cdots\cup C_{k}$ we mean the chain of cycles obtained from $\Gamma$ by deleting all cycles $C_{k+1}\cup\cdots\cup C_{g}$ and the segments from $w_{i}$ to $v_{i+1}$ for $k\leq i\leq g-1$. Its torsion profile is given by $(m_{2},\cdots,m_{k})$; we still talk about $\underline{m}$-displacement tableaux on $C_{1}\cup\cdots\cup C_{k}$ (so we use the same notation to indicate the torsion profile).
Let $g^{\prime}=t(g-d+r,r)$, hence $g^{\prime}\leq g-1$. Then $t$ restricted to $[(g-d+r)\times r]=(g^{\prime}-(d-(g-g^{\prime})-1)+(r-1))\times((r-1)+1)]$ is an $\underline{m}$-displacement tableau for $C_{1}\cup\cdots\cup C_{g^{\prime}}$. Since $r\geq 2$ of course $r-1\geq 1$ while $g^{\prime}-(d-(g-g^{\prime})-1)+(r-1)=g-d+r\geq 2$. From the induction hypothesis we obtain an $\underline{m}$-displacement tableau $\tilde{t^{\prime}}$ on $[(g^{\prime}-(d-(g-g^{\prime})-1)+2(r-1)-1)\times 2]=[(g-d+2r-2)\times 2]$ for $C_{1}\cup\cdots\cup C_{g^{\prime}}$. In case $g^{\prime}\leq g-2$ then we enlarge $\tilde{t^{\prime}}$ with $t^{\prime}(g-d+2r-1,1)=g-1$ and $t^{\prime}(g-d+2r-1,2)=g$ and we obtain an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ for $\Gamma$.
To continue the proof we can assume that $g^{\prime}=t(g-d+r,r)=g-1$, hence also $(g-d+r,r+1)=g$. In case $t(g-d+r-1,r+1)=g-1$ then $m_{g-1}=2$ and by extending $\tilde{t^{\prime}}$ with $t^{\prime}(g-d+2r-1,1)=g-1$ and $t^{\prime}(g-d+2r-1,2)=g$ we obtain an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ for $\Gamma$.
To continue the proof we can assume that $t(g-d+r-1,r+1)\leq g-2$. It implies $g^{\prime}=t(g-d+r-1,r)\leq g-3$ (so we changed the meaning of the symbol $g^{\prime}$), so the restriction of $t$ to $[(g-d+r-1)\times r]=[(g^{\prime}-(d-(g-g^{\prime}))+(r-1)\times((r-1)+1)]$ is an $\underline{m}$-displacement tableau for $C_{1}\cup\cdots\cup C_{g^{\prime}}$ with $r-1\geq 1$ and $g^{\prime}-(d-(g-g^{\prime}))+(r-1)=g-d+r-1$. In case $g-d+r=2$ then from $d\leq g-1$ we obtain $r\leq 1$ contradicting our assumption $r\geq 2$. So we find $g-d+r-1\geq 2$ and we can use the induction hypothesis giving the existence of an $\underline{m}$-displacement tableau $\tilde{t^{\prime}}$ on $[(g^{\prime}-(d-(g-g^{\prime}))+2(r-1)-1)\times 2]=[(g-d+2r-3)\times 2]$ on $C_{1}\cup\cdots\cup C_{g^{\prime}}$. In case $g^{\prime}\leq g-4$ we can extend $\tilde{t^{\prime}}$ putting $t^{\prime}(g-d+2r-2,1)=g-3$, $t^{\prime}(g-d+2r-2,2)=g-2$, $t^{\prime}(g-d+2r-1,1)=g-1$ and $t^{\prime}(g-d+2r-1,2)=g$. We obtain an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ for $\Gamma$.
To continue the proof we can assume $t(g-d+r-1,r)=g-3$ and hence also $t(g-d+r-1,r+1)=g-2$.
Now we consider the following claim involving a positive integer $k$.
Claim.
Let $k$ be an integer satisfying $11\leq k\leq g-d+r-$ 1. If there exist some $0\leq k^{\prime}\leq k$ such that $t(g-d+r-k^{\prime},r+1)\neq g-2k^{\prime}$ or $t(g-d+r-k^{\prime},r)\neq g-2k^{\prime}-1$ then $\Gamma$ has an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$.
In case Theorem 2 would not be true then the claim implies $g-2k\geq g-d+2r-k$, i.e. $k\leq d-2r$. Also in case $k=d-2r$ it would imply $t(i,j)=i+j-1$ for $1\leq i\leq g-d+r-k$ and $11\leq j\leq r+$ 1. In the above arguments we proved the claim for $k=1$. In case $k<d-2r$ and $k<g-d+r-1$ we are going to prove that the claim implies the similar claim replacing $k$ by $k+1$. In the other cases we are going to prove that the claim implies the existence of an $\underline{m}$-replacement tableau on $[(g-d+2r-1)\times 2]$. In this way the proof of Theorem 2 becomes complete.
First we assume the claim holds for some value $k$ satisfying $k<d-2r$ and $g-d+r-k>1$ and we assume no value $k^{\prime}$ as in the claim exists. Restricting $t$ to $[(g-d+r-k)\times r]=[((g-2k-1)-(d-k-2)+(r-1))\times((r-1)+1)]$ we obtain an $\underline{m}$-displacement tableau on $C_{1}\cup\cdots\cup C_{g-2k-1}$. We have $r-1\geq 1$ and $(g-2k-1)-(d-k-2)+(r-1)=g-k-d+r\geq 2$. By assumption we obtain an $\underline{m}$-tableau $\tilde{t^{\prime}}$ on $[((g-2k-1)-(d-k-2)+2(r-1)-1)\times 2]=[(g-d+2r-k-2)\times 2]$ for $C_{1}\cup\cdots\cup C_{g-2k-1}$. In case $t(g-d+r-k-1,r+1)=g-2k-1$ we have $m_{g-2k-1}=2$. Then we can extend $\tilde{t^{\prime}}$ by $t^{\prime}(g-d+2r-k-2+x,1)=g-2k-3+2x$ and $t^{\prime}(g-d+2r-k-2+x,2)=g-2k-2+2x$ for $11\leq x\leq k+$ 1, which gives rise to an $\underline{m}$-displacement tableau of $\Gamma$.
To continue the proof we can assume $t(g-d+r-k-1,r+1)\leq g-2k-2$. In case $t(g-d+r-k-1,r)\leq g-2k-4$ then the restriction of $t$ to $[(g-d+r-k-1)\times r]=[((g-2k-4)-(d-k-4)+(r-1))\times((r-1)+1)]$ is an $\underline{m}$-displacement tableau for $C_{1}\cup\cdots\cup C_{g-2k-4}$. We have $r-1\geq 1$ and $(g-2k-4)-(d-k-4)+(r-1)=g-d-k+r-1\geq 1$.
In case $g-d-k+r=2$ then the restriction of $t$ to $[(g-d+r-k)\times(r+1)]=[((g-2k)-(d-k)+r)\times(r+1)]$ is an $\underline{m}$-displacement tableau on $C_{1}\cup\cdots\cup C_{g-2k}$. We have $(g-2k)-(d-k)+r=g-d-k+r\geq 2$. In case $d-k\leq(g-2k)-1$ we have $g\geq d+k+1$, so $d+k+2-r\geq d+k+1$ and therefore $r\leq 1$, contradicting our assumption. So we find $d-k\geq g-2k$ and then Lemma 1 implies the existence of an $\underline{m}$-displacement tableau $\tilde{t^{\prime}}$ on $[((g-2k)-(d-k)+2r-1)\times 2]=[(g-k-d+2r-1)\times 2]$ for $C_{1}\cup\cdots\cup C_{g-2k}$. Using the remaining $2k$ values $\{g-2k+1,\cdots,g\}$ we can complete $\tilde{t^{\prime}}$ to an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ for $\Gamma$.
In case $g-d-k+r\geq 3$ then by assumption we find an $\underline{m}$-displacement tableau $\tilde{t^{\prime}}$ on $[((g-2k-4)-(d-k-4)+2(r-1)-1)\times 2]=[(g-k-d+2r-3)\times 2]$ on $C_{1}\cup\cdots\cup C_{g-2k-4}$. We can extend $\tilde{t^{\prime}}$ by $t^{\prime}(g-k-d+2r-3+i,1)=g-2k-4+2i-1$ and $t^{\prime}(g-k-d+2r-3+i,2)=g-2k-4+2i$ for $21\leq i\leq k+$ 2 and we obtain an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ on $\Gamma$.
To continu the proof we can assume $t(g-d+r-k-1,r)=g-2k-3$ and therefore also $t(g-d+r-k-1,r+1)=g-2k-2$. So we obtain the claim with $k$ replaced by $k+1$.
Next we assume the claim holds for some value $k$ with $g-d+r-k>1$ but $k=d-2r$. As remarked we have $t(i,j)=i+j-1$ for $1\leq i\leq g-d+r-k$ and $11\leq j\leq r+$ 1. It follows $m_{i}=2$ for $12\leq i\leq g-d+2r-k-$ 1. So $\tilde{t^{\prime}}(i,j)=i+j-1$ is an $\underline{m}$-displacement tableau on $[(g-d+2r-k-1)\times 2]$ for $C_{1}\cup\cdots\cup C_{g-d+2r-k}$. But the $d+k-2r=2k$ values from $\{g-d+2r-k+1,\cdots,g\}$ can be used to extend $\tilde{t^{\prime}}$ to an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ for $\Gamma$.
Finally we assume $g-d+r-k=1$. Then the restriction of $t$ to $[2\times(r+1)]=[((g-2k+2)-(g+r-2k)+r)\times(r+1)]$ is an $\underline{m}$-displacement tableau on $C_{1}\cup\cdots\cup C_{g-2k+2}$. Note that $g+r-2k\geq 0$ since $g-2k\geq 0$. Also $(g-2k+2)-(g+r-2k)+r=2$ but $r+g-2k\geq g-2k+2$. So from Lemma 1 we obtain there exists an $\underline{m}$-displacement tableau $\tilde{t^{\prime}}$ on $[((g-2k+2)-(g+r-2k)+2r-1)\times 2]=[(r+1)\times 2]$ for $C_{1}\cup\cdots\cup C_{g-2k+2}$. The set $\{g-2k+3,\cdots,g\}$ consists of $22k-$ 2 unused values in $\tilde{t^{\prime}}$. Since $(r+1)+(k-1)=(d-r+k+1)-d+2r-1=g-d+2r-1$ those values can be used to obtain an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ for $\Gamma$.
This finishes the proof of Theorem 2, hence also of Theorem 1.
{bibsection}
BakerM.JensenD.Degeneration of linear series from the tropical point of view and applicationsM. Baker, S.Payne (eds) Nonarchimedean and Tropical GeometrySimons Symposia2016365–433@article{ref1, author = {M. Baker}, author = {D. Jensen}, title = {Degeneration of linear series from the tropical point of view and applications}, journal = {M. Baker, S.Payne (eds) Nonarchimedean and Tropical Geometry}, series = {Simons Symposia}, year = {2016}, pages = {365-433}} Cook-PowellK.JensenD.Tropical methods in hurwitz-brill-noether theoryAdvances in Math.3982022108199@article{ref5, author = {K. Cook-Powell}, author = {D. Jensen}, title = {Tropical methods in Hurwitz-Brill-Noether theory}, journal = {Advances in Math.}, volume = {398}, year = {2022}, pages = {108199}} CoolsF.DraismaJ.PayneS.RobevaE.A tropical proof of the brill-noether theoremAdvances in Math.2302012759–776@article{ref4, author = {F. Cools}, author = {J. Draisma}, author = {S. Payne}, author = {E. Robeva}, title = {A tropical proof of the Brill-Noether Theorem}, journal = {Advances in Math.}, volume = {230}, year = {2012}, pages = {759-776}} CoolsF.PanizzutM.The gonality sequence of complete graphsElectr. J. Comb.24(4)2017P4.1@article{ref6, author = {F. Cools}, author = {M. Panizzut}, title = {The gonality sequence of complete graphs}, journal = {Electr. J. Comb.}, volume = {24(4)}, year = {2017}, pages = {P4.1}} CoppensM.MartensG.Secant spaces and clifford’s theoremComposition Math.781991193–212@article{ref2, author = {M. Coppens}, author = {G. Martens}, title = {Secant spaces and Clifford's Theorem}, journal = {Composition Math.}, volume = {78}, year = {1991}, pages = {193-212}} DeveauA.JensenD.KainicJ.MitropolskyD.Gonality of random graphsInvolve92016715–720@article{ref3, author = {A. Deveau}, author = {D. Jensen}, author = {J. Kainic}, author = {D. Mitropolsky}, title = {Gonality of random graphs}, journal = {Involve}, volume = {9}, year = {2016}, pages = {715-720}} PfluegerN.Special divisors on marked chains of cyclesJ. Combin. Theory Ser. A1502017182–207@article{ref7, author = {N. Pflueger}, title = {Special divisors on marked chains of cycles}, journal = {J. Combin. Theory Ser. A}, volume = {150}, year = {2017}, pages = {182-207}}